# INQUIRE-Search: A Framework for Interactive Discovery in Large-Scale Biodiversity Databases

### v3

@@ 2.3.2 CS2: Post-fire Forest Regrowth
Overview: Understanding how forests recover following fire is essential for assessing long-term ecosystem resilience, particularly in the western United States where wildfire has been the dominant disturbance for millennia [whitlock2008long, marlon2012long] and many forests are fire-adapted [keeley2011fire]. Recent warming and drying trends associated with climate change, coupled with the legacy of fire suppression, are driving fires that are more frequent, severe, and intense [abatzoglou2016impact, juang2022rapid], exceeding the adaptive capacity of many forest systems by limiting post-fire regeneration and altering species composition [harvey2016burn, davis2023reduced]. Despite the ecological importance of post-fire recovery for understanding forest resilience, successional trajectories, and long-term ecosystem change, existing approaches provide limited capacity to assess fine-scale regrowth: satellite data often lack the spatial or spectral resolution needed to detect young trees or differentiate species [xu2021spatial, kiel2022trees], while field surveys are constrained by cost and limited spatial coverage. Because community-sourced photographs can reveal early regeneration that is invisible to both satellites and sparse field campaigns, this case study evaluates whether INQUIRE-Search can identify young coniferous and deciduous individuals within the 2012 High Park Fire perimeter and examine how their occurrence varies across burn severity classes (Fig. 3).
Search Strategy: We targeted two functional groups (young coniferous trees and young deciduous trees), using text queries such as "young coniferous trees in burned forest" and "young deciduous trees in burned forest". Spatial filters restricted observations to the High Park Fire perimeter: latitude (40.57 to 40.75 N) and longitude (105.18 to 105.54 W) (Fig 3). Temporal filters restricted images to post-fire dates. After retrieval, we applied a finer spatial filter using reported image coordinates, retaining only images within the High Park MTBS fire perimeter via the sf package in R3 [pebesma2018simple].
Verification: An image was marked informative when it contained the forest community of interest (i.e., an individual of the target forest type could be clearly distinguished). We allowed individual images to be marked as informative in multiple searches if the image contained individuals representing both forest types (e.g. a coniferous and deciduous tree seedling). While processing images for searches related to trees, we also decided to include trees and shrubs within the same target forest type as it was difficult to distinguish young trees from shrubs in images. Top 200 images were inspected.
Analysis: Coordinates of informative images were combined with MTBS burn severity categories by aggregating the 30 m burn data to the 0.01 precision typical of iNaturalist coordinates and assigning each location the modal severity class using the terra package in R4 [hijmans2022package]. We then counted marked images by forest type and burn severity to evaluate how regeneration outcomes varied across burn categories.
2.3.3 CS3: Wildlife Mortality
@@
Overview: Understanding patterns of avian mortality is fundamental to conservation efforts, particularly given the loss of nearly three billion birds in North America since 1970 [rosenberg2019decline]. Mortality risk varies seasonally as birds encounter different threats across the annual cycle [loss2015direct, marra2015call], and collisions with human-made structures–one of the leading anthropogenic causes of bird death [loss2014estimation]–show strong seasonal peaks during migration [riding2021multi, scott2023causes] and vary across landscapes due to differing local hazards and bird abundances [hager2017continent]. Yet documenting spatiotemporal patterns of mortality is challenging because mortality events are rare, and alternative monitoring methods such as marking or tracking individuals are resource-intensive and limited in scale [yanco2025tracking]. Since large-scale image repositories could reveal where and when birds die, this case study uses INQUIRE-Search to investigate seasonal mortality dynamics in an urban (Boston, MA) and rural (Pioneer Valley, MA) setting and compare relative mortality between these contrasting anthropogenic contexts (Fig 4).
Search Strategy: We queried for "dead bird" filtering by each of the two equal-sized bounding boxes using location filters. We use latitude (42.31 to 42.38 N) and longitude (-71.14 to -71.01 W) for the urban study area, and latitude (42.31 to 42.38 N) and longitude (-72.64 to -72.51 W) for the rural study area.
Verification: We excluded images of live birds, single features, non-avian taxa, and duplicate observations of the same event (duplicate sharing species, month, latitude and longitude) to identify true avian mortality events.
Analysis: To control for differing bird abundance, observation effort, and mortality among sites, we calculate the mortality index defined as:
@@
Here, $D_{m,s}$ is the mortality count and $O_{m,s}$ is the observation count for month $m$ at site $s$. Values above zero indicate fold increases in mortality relative to the site’s annual mean, while values below zero indicate fold decreases. The total observation count was extracted using the rinat package in R [barve2022rinat].
2.3.4 CS4: Resolving Plant Phenology
Overview: Studying the timing of life-history events is a primary way researchers detect biodiversity responses to global change [parmesan2003globally], as phenological shifts in processes such as germination, flowering, and senescence can serve as early warning signals that precede demographic, genetic, or geographic change. While fine-scale phenological datasets collected through repeated field surveys provide valuable local insight [austin2024climate], detecting and attributing phenological change at broader spatial scales remains challenging due to limited temporal resolution and data availability [doi2017macroecological]. Although some macrophenological studies have used computer vision to extract phenological information from digitized specimens [williamson2025long], most rely on coarse metrics such as mean flowering date because intra-annual data on distinct phenophases are sparse. The growing volume of geotagged plant observations on community-science platforms like iNaturalist offers a path to overcoming these limitations, provided phenological stages can be systematically identified from photographs. In this case study, we test whether INQUIRE-Search can recover detailed phenophases—emergence, flowering, seeding, and senescence—for common milkweed (Asclepias syriaca) in southern Québec (Fig 5).
@@ 2.3.4 CS4: Resolving Plant Phenology
Search Strategy: We created stage-specific text queries targeting (1) emergence ("Milkweed germinating or emerging"), (2) flowering ("Milkweed flowering"), (3) seeding ("Milkweed producing seeds or milkweed with seeds"), and (4) senescence ("Milkweed dying or withering or senescence"). We also used a species filter ("Asclepias syriaca") and a geographic filters for latitude (45.03 to 46.54 N) and longitude (-74.68 to -71.66 W) to limit observations to southern Quebec.
Verification: Top 200 images were labeled as informative according to strict morphological criteria:
  • Emergence: <4<4 pairs of adult leaves
  • Flowering: presence of open petals
  • Seeding: open pods with visible seeds
  • Senescence: absence of green leaves
Analysis: Observation dates were converted to Day-of-Year (DOY). Mean DOY values were compared among stages using ANOVA, followed by Tukey’s HSD to identify pairwise differences.
2.3.5 CS5: Whale Re-Identification
Overview: Animal re-identification – the ability to recognize individual animals across space and time – is essential for estimating population size and survival, reconstructing migratory routes, evaluating site fidelity, and monitoring long-term shifts in habitat use across different species and ecosystems [krebs1989ecological]. For wide-ranging marine mammals such as humpback whales, these insights depend on longitudinal photo-ID datasets linking repeated sightings across years and ocean basins [katona1981identifying, calambokidis2001movements, martin1984migration], typically using pigmentation and trailing-edge patterns on the ventral fluke [howard2018humpback]. Although many fluke images exist in open repositories like iNaturalist, scaling re-ID remains constrained by the difficulty of locating images that show a clear, properly oriented fluke suitable for identification, as existing filters cannot reliably distinguish low-quality, mis-angled, or incomplete views (Fig 6) [otarashvili2024multispecies]. Because identifiable fluke photographs are scattered across diverse image sources with variable pose, lighting, and metadata, this case study evaluates whether INQUIRE-Search can surface individually identifiable humpback whale flukes from unstructured community-science imagery and match them to known individuals in the HappyWhale dataset.
Figure 6: (Left) Re-Identification of humpback whales. Unique Re-Identification of individuals is possible through careful examination of the fluke. (Right) iNaturalist humpback whale images with "fluke" description/tag filtering. Recent iNaturalist humpback whale images, even with the fluke filtering, are not identifiable due to the photographed angle or image quality.
Search Strategy: We queried for "white underside of humpback whale fluke" with the humpback whale species filter.
Verification: Images were considered informative only if the ventral fluke was clearly visible, unobstructed, and in focus. Top 200 retrievals were inspected.
Analysis: For analysis, we assess whether any iNaturalist retrievals corresponded to known individuals in the HappyWhale dataset. To accelerate matching, we cropped each image using Grounding DINO [liu2024grounding] and embedded both the iNaturalist and HappyWhale datasets using a multi-species re-identification model [otarashvili2024multispecies]. For each of the verified INQUIRE-Search images, we retrieved the top three closest HappyWhale candidates by embedding similarity and manually reviewed these candidates to confirm matches. Although not conducted by experts, humpback flukes possess distinct and easily recognizable pigmentation and trailing-edge patterns that allow reliable identification by non-experts.
3 Results
3.1 CS1: Seasonal Variation in Bird Diets
Results and Analysis
Our analysis revealed substantial variation in the availability of dietary information across the five study species shown in Fig 7. American Robin yielded the highest number of informative (i.e., containing dietary information) images (669 total), followed by Red-bellied Woodpecker and American Tree Sparrow (395 and 188, respectively), while for both Ancient Murrelet and Gray-cheeked Thrush we found a negligible number of images containing dietary information. Among the species for which dietary information was available, seasonal patterns in image availability were also evident: American Robin had more informative images in summer than winter, while Red-bellied Woodpecker and American Tree Sparrow had more images available for winter months.
Figure 7: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for "Red-bellied Woodpecker" and "Invertebrate." (Bottom) Diet comparisons. INQUIRE-Search yields dietary patterns that align closely with those documented in the SAviTraits reference dataset.
For the three species with sufficient data, dietary composition derived from INQUIRE-Search closely matched reference data from SAviTraits (Fig 7). For example, in summer, American Robin exhibited an even split between fruit and invertebrate consumption (50% each) in both datasets, while Red-bellied Woodpecker showed a diet dominated by seeds (50% in INQUIRE-Search vs. 40% in SAviTraits), with smaller contributions from invertebrates, fruits, and plant material. In winter, all three species exhibited diets strongly dominated by a single food type, and we again observed close agreement between INQUIRE-Search outputs and SAviTraits data.
Findings
We demonstrate that our method can effectively reproduce dietary patterns documented in established literature, provided that sufficient image data are available. The close alignment between INQUIRE-Search-derived avian diets and SAviTraits reference data highlights the potential of community science imagery to yield accurate ecological insights when properly analyzed through a systematic search methodology. However, this approach is most effective for commonly photographed species such as American Robin and Red-bellied Woodpecker. Two of the species we examined lacked sufficient image data for meaningful analysis despite being well-studied in the scientific literature and represented in databases like SAviTraits, underscoring a fundamental constraint of community-sourced data. This limitation is likely even more pronounced for understudied species—precisely those for which new dietary data are most needed—because they are also underrepresented in iNaturalist and thus rarely retrievable with INQUIRE-Search.
3.2 CS2: Post-fire Forest Regrowth
Figure 8: (Top) Query & Verification. Filtering top INQUIRE-Search returns for "young coniferous trees in burned forest." (Bottom) Tree observations across burn severity. Distribution of observed young coniferous and deciduous trees in community science collected images show strong relationships with burn severity regions.
Results and Analysis
Of the 200 images exported from INQUIRE-Search for each search, there were 100 (coniferous) and 112 (deciduous) images with reported coordinates within the High Park burn perimeter. We marked more informative images for deciduous forest species than coniferous (78 vs. 45 images, respectively). We also marked five images depicting both coniferous and deciduous individuals, indicating areas that could recover into mixed forest communities.
Among images marked for analysis, we found strong relationships between burn severity and post-fire regeneration of both coniferous and deciduous forest types. For both forest types, most marked images of regenerating vegetation occurred in areas classified as unburned-to-low or low burn severity, with few observations in moderate or high severity areas (Fig 8, left). However, coniferous regeneration showed a stronger negative association with increasing burn severity: 93.3% of coniferous images were located in unburned-to-low or low severity areas, with only 4.4% and 2.2% in moderate and high burn severity areas, respectively (Fig 8, right). In contrast, deciduous regeneration was somewhat more evenly distributed, with 84.6% of marked images in unburned-to-low or low severity areas, and a higher proportion (7.6% each) in areas burned at moderate and high severity (Fig 8, right). These patterns suggest that post-fire forest regeneration, as captured by community-science images, is more common in lower severity areas, with coniferous species showing greater sensitivity than deciduous species.
Findings
In this case study, we demonstrate that INQUIRE-Search can be used to effectively identify post-fire forest regeneration patterns across burn severity gradients. Our approach successfully distinguished between regenerating coniferous and deciduous forest types and provided spatially explicit insights into post-disturbance recovery within the first decade after the High Park fire. Our findings are consistent with previous research documenting that increased fire severity decreases conifer regeneration likelihood both regionally and specifically after the High Park fire [davis2023reduced, wright2017post].
Several limitations remain. Image coordinates are rounded to 0.01 to match the limiting spatial precision of many iNaturalist records, but this prevents fine-scale matching to burn severity or site attributes such as seed source or soil moisture. Classification from user-submitted photographs also introduces uncertainty, as landscape images often obscure species identification, and deciduous seedlings, saplings, shrubs, and herbaceous plants can appear morphologically similar. Finally, iNaturalist data are opportunistic and biased toward accessible, aesthetically appealing areas, which may result in severely burned sites being underrepresented (Fig 3). Thus, while INQUIRE is a powerful supplemental tool for assessing post-fire recovery, its results should be interpreted alongside field data and remote-sensing derived products to minimize bias.
3.3 CS3: Wildlife Mortality
Results and Analysis
The results show distinct patterns between urban and rural settings. Boston exhibits notable peaks in mortality during the spring and fall, moderate mortality rates during the summer, and low mortality rates in winter. In contrast, the Pioneer Valley shows elevated mortality rates in summer and fall, moderate mortality rates in early winter, and low mortality rates in late winter and spring. The high mortality rates at both sites in September and October are consistent with expectations of peak migratory passage—and the concomitant risk of building collisions—being highest in the fall in the eastern United States [horton2019bright]. Interestingly, Boston exhibited higher mortality rates during spring migration which may reflect differential migratory passage rates or other differences in site-specific risk factors [scott2023causes]. Finally, the Pioneer Valley did not exhibit the winter drop in mortality rates that we observed in Boston. While this could reflect differences in the avian communities or risks between the two sites, we cannot rule out the possible role of increased detection probability due to reduced vegetative cover disproportionately affecting the more rural Pioneer Valley.
Findings
We demonstrate that INQUIRE-Search can detect distinct spatiotemporal patterns in avian mortality through analysis of community-sourced photographs. The contrasting seasonal dynamics between urban (Boston) and rural (Pioneer Valley) sites provide a compelling demonstration of the importance of anthropogenic context in driving avian demographics.
However, several important limitations must be considered when interpreting these results. First, detection biases differ substantially across landscapes. Urban carcasses occur in high-visibility areas and are readily detected, while rural mortality events are often missed because they occur in remote or forested habitats where scavenging and decomposition quickly eliminate evidence. Additionally, our mortality index assumes that the ratio of dead bird observations to total bird observations provides a meaningful proxy for actual mortality rates, but this assumption may not hold if detection probabilities vary systematically across seasons, locations, or species.
3.4 CS4: Resolving Plant Phenology
@@
Figure 10: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for "flowering" phenological stage. (Bottom) Phenological stages occur at significant times throughout the year. (left) density of observations from INQUIRE-Search corresponding to different stages. (right) distribution of observations from INQUIRE-Search in different phenological stages visualized with results from ANOVA/Tukey test.
Results and Analysis
Our analysis shows significant differences in timing for each phenological stage resolved using INQUIRE-Search for common milkweed in southern Quebec. The phenological stage with the best return rate (number of observations matching the queried stage) was flowering (84.5%) followed by seeding (80.5%), senescence (26%), and finally emergence (22.5%). An ANOVA revealed significant differences in stages for the day of year of returned observations, and pairwise differences between stages was tested using a Tukey test. All stages were significantly (p<0.05) different from one another with the exception of emergence and flowering, which showed considerable overlap in timing.
Findings
This case study shows that INQUIRE-Search can be used systematically to resolve detailed phenological information at regional scales for specific species. Some phenological stages, namely, flowering and seeding, were more successfully resolved than emergence and senescence. While this could be a limitation of the models ability to identify relevant observations to those stages (the difference between an emerging plant and a young plant is minor) it is also likely due to bias in the community science observations queried, as observations of flowers and seeds are likely more common as they can aid taxonomic identification [iNaturalist202xHighQualityObservations]. Resolving patterns of senescence will also require specific considerations to account for the fact that senescent plants can also be observed long after senescence, evidenced by the fact we identified observations of senescent milkweed throughout the year. Nonetheless, this sort of fine-scale phenological information would have traditionally required daily or weekly surveys to resolve at the level of a single site, let alone a region like Southern Quebec. Our results demonstrate that INQUIRE-Search has the potential to drastically increase our ability to extract phenological information from the wealth of rapidly growing community science observations.
3.5 C5: Whale Re-Identification
Results and Analysis

### v5

@@ 3.2 CS2: Post-fire forest regrowth
Recent warming and drying trends associated with climate change are driving wildfires that are more frequent, severe, and intense (Abatzoglou and Williams, 2016; Juang et al., 2022), exceeding the adaptive capacity of many forest systems (Harvey et al., 2016; Davis et al., 2023). Understanding post-fire recovery is essential to assessing long-term forest resilience, successional trajectories, and long-term ecosystem change (Whitlock et al., 2008; Marlon et al., 2012; Keeley et al., 2011). However, existing approaches limit the study of fine-scale regrowth: satellite data often lack the spatial or spectral resolution needed to detect young trees or differentiate species (Xu et al., 2021; Kiel and Turner, 2022), while field surveys are constrained by cost and limited spatial coverage. In contrast, community-sourced photographs can reveal early regeneration that is invisible to both satellites and sparse field campaigns.
INQUIRE-Search workflow for CS2. 1) Query and Prioritize: We targeted two functional groups (young coniferous trees and young deciduous trees), using text queries such as “young coniferous trees in burned forest” and “young deciduous trees in burned forest.” Spatial filters restricted observations to the High Park Fire perimeter: latitude (40.57 to 40.75 N) and longitude (105.18 to 105.54 W) (Figure 4). Temporal filters restricted images to post-fire dates. After retrieval, we applied a finer spatial filter using reported image coordinates, retaining only images within the High Park MTBS fire perimeter via the sf package in R3 (Pebesma, 2018). 2) Retrieve and Verify: From the top 200 images retrieved for each functional group, we marked images that clearly contained at least one individual of the target forest type. We allowed individual images to be marked as informative in multiple searches if the image contained individuals representing both forest types (e.g. a coniferous and deciduous tree seedling). We also decided to include trees and shrubs within the same target forest type as it was difficult to distinguish young trees from shrubs in images. 3) Export and Analyze: Coordinates of informative images were combined with MTBS burn severity categories by aggregating the 30 m burn data to the 0.01 precision typical of iNaturalist coordinates and assigning each location the modal severity class using the terra package in R4 (Hijmans et al., 2022). We then counted marked images by forest type and burn severity to evaluate how regeneration outcomes varied across burn categories.
Results. INQUIRE-Search successfully identified photographs depicting early recovery within the High Park Fire boundary (Table 3). After fine-grained spatial filtering, we reviewed $N_{insp}=112$ deciduous and $N_{insp}=100$ coniferous images, verifying $N_{ret}=78$ deciduous ($Y=0.70$) and $N_{ret}=45$ coniferous ($Y=0.45$) regeneration events.
Without INQUIRE-Search, baseline retrieval relied on geographic and temporal filtering, followed by inspection of the first $N_{insp}=200$ images within the fire boundary. This yielded only $N_{ret}=6;Y=0.03$ coniferous and $N_{ref}=19;Y=0.095$ deciduous shrubs or seedlings. Under comparable manual inspection budgets, INQUIRE-Search produced substantially denser datasets, enabling subsequent analysis of regeneration patterns across burn-severity gradients.
Figure 5: (Top) Query & Verification. Filtering top INQUIRE-Search returns for “young coniferous trees in burned forest.” (Bottom) Tree observations across burn severity. Distribution of observed young coniferous and deciduous trees in community science collected images show strong relationships with burn severity regions.
Using the curated INQUIRE-Search dataset, we found strong relationships between burn severity and post-fire regeneration. Most regeneration occurred unburned-to-low or low burn severity areas, with few observations in moderate or high severity areas (Figure 5, bottom left). However, the strength of this pattern differed between functional groups. Coniferous regeneration showed a stronger negative association with increasing burn severity, while deciduous regeneration was somewhat more evenly distributed (Figure 5, bottom right). These patterns are consistent with previous research documenting that increased fire severity decreases conifer regeneration likelihood following the High Park fire and across regional forests(Davis et al., 2023; Wright and Rocca, 2017).
Several limitations remain, including coarse coordinate precision (images coordinates are rounded to 0.01 to match the limiting spatial precision of many iNaturalist records), challenges distinguishing seedlings and shrubs in photographs, and biases toward accessible locations, which may underrepresent severely burned patches (Figure 4).
3.3 CS3: Wildlife mortality
Overview. This case study explores the use of INQUIRE-Search to collect mortality instances in urban (Boston, MA) and rural (Pioneer Valley, MA) settings across time to investigate seasonal mortality dynamics and to compare relative mortality across different anthropogenic contexts (Figure 6).
@@ 3.3 CS3: Wildlife mortality
Understanding avian mortality is central to conservation, particularly given the loss of nearly three billion birds in North America since 1970 (Rosenberg et al., 2019). Mortality risk varies seasonally (Loss et al., 2015; Marra et al., 2015) and is strongly influenced by collisions with human-made structures such as windows and vehicles(Loss et al., 2014), which peak during migration (Riding et al., 2021; Scott et al., 2023) and differ across landscapes due to differing landscapes and bird abundances (Hager et al., 2017). Yet documenting spatiotemporal mortality patterns is challenging because events are rare, and alternative monitoring methods such as marking or tracking individuals are resource-intensive and limited in scale (Yanco et al., 2025). Since community science observations are often collected in close proximity to human activity, they frequently capture wildlife mortality events that would otherwise go unrecorded.
INQUIRE-Search workflow for CS3. 1) Query and Prioritize: We queried for “dead bird” within two equal-sized bounding boxes using location filters: an urban site (Boston) (42.31 to 42.38 N, -71.14 to -71.01 W) and a rural site (Pioneer Valley) (42.31 to 42.38 N, -72.64 to -72.51 W). 2) Retrieve and Verify: To identify true avian mortality events, we excluded images of live birds, single anatomical features, non-avian taxa, and duplicate observations of the same event (defined as records sharing species, month, latitude, and longitude). Images were reviewed sequentially in ranked order, and verification terminated after 200 consecutive non-relevant retrievals. Using this criterion, we verified $N_{insp}=359$ images in Pioneer Valley. In Boston, where upload density was substantially higher, we capped verification at $N_{insp}=1000$ retrievals due to a limited annotation budget, although additional relevant observations may have been present beyond this threshold. 3) Export and Analyze: To control for differing bird abundance, observation effort, and mortality among sites, we calculate the mortality index defined as:
@@
Here, $D_{m,s}$ is the mortality count and $O_{m,s}$ is the observation count for month $m$ at site $s$. Values above zero indicate an increase in mortality relative to the site’s annual mean, while values below zero indicate a decrease. The total observation count was extracted using the rinat package in R (Barve and Hart, 2022).
Results. Table 4 show that INQUIRE-Search surfaced hundreds of verified avian mortality events in both rural and urban settings. In Boston, inspection of $N_{insp}=1000$ yielded $N_{ret}=545$ and in Pioneer Valley, $N_{ret}=79$ events were verified from $N_{insp}=360$ images.
In comparison, existing iNaturalist filters recovered significantly fewer mortality events. iNaturalist includes structured annotations for “dead/alive” attributes that observers can optionally provide. However, applying this filter yielded only $N_{ret}=295$ dead-bird observations in Boston and $N_{ret}=42$ in Pioneer Valley. Using keyword queries performed even worse, returning just $N_{ret}=46$ and $N_{ret}=8$ candidate images in Boston and Pioneer Valley respectively. In many cases, baseline searches returned fewer total images than the nominal inspection budget, indicating that mortality events are sparsely indexed in structured metadata and text fields and are difficult to recover using conventional metadata- and keyword-based search alone.
Using the verified datasets, we examined seasonal mortality dynamics across the two landscapes (Figure 7). Both sites exhibited high mortality during fall migration (September-October), consistent with increased collision risk in the eastern United States (Horton et al., 2019). Interestingly, Boston exhibited higher mortality rates during spring migration, potentially reflecting differences in migratory passage rates or site-specific risk factors (Scott et al., 2023). In contrast, Pioneer Valley lacked a pronounced winter decline, which may reflect differences in community composition or potentially increased winter detectability due to reduced vegetation, which would disproportionally affect rural data collection.
Overall, these findings show that INQUIRE-Search can detect distinct spatiotemporal patterns in avian mortality that are difficult to capture using metadata alone. The contrasting seasonal dynamics between urban and rural sites highlight the role of anthropogenic context in driving avian demographics. However, interpretation must account for detection biases: urban carcasses are more visible, while rural mortality may be underrepresented due to scavenging and remoteness. Additionally, our mortality index assumes that the ratio of dead bird observations to total bird observations provides a meaningful proxy for actual mortality rates, but this assumption may not hold if detection probabilities vary systematically across seasons, locations, or species.
3.4 CS4: Resolving plant phenology
Overview. In this case study, we test whether INQUIRE-Search can recover four distinct phenophases—emergence, flowering, seeding, and senescence—for common milkweed (Asclepias syriaca) in southern Québec from community science images (Figure 8).
@@ 3.4 CS4: Resolving plant phenology
Studying the timing of life-history events is a primary way researchers detect biodiversity responses to global change (Parmesan and Yohe, 2003). While fine-scale phenological datasets from repeated field surveys provide local insight (Austin et al., 2024), phenological change remains difficult to assess at broader spatial scales (Doi et al., 2017). Although some existing studies have used computer vision to extract phenological information (Williamson et al., 2025), most rely on coarse metrics such as mean flowering date because data on distinct phenophases are sparse. The growing volume of plant observations on community-science platforms like iNaturalist offers a path to overcoming these limitations, provided phenological stages can be identified from photographs.
INQUIRE-Search workflow for CS4. 1) Query and Prioritize: We created stage-specific text queries targeting each phase: (1) emergence (“Milkweed germinating or emerging”), (2) flowering (“Milkweed flowering”), (3) seeding (“Milkweed producing seeds or milkweed with seeds”), and (4) senescence (“Milkweed dying or withering or senescence”). We also used a species filter (“Asclepias syriaca”) and a geographic filters for latitude (45.03to 46.54 N) and longitude (-74.68 to -71.66 W) to limit observations to southern Quebec. 2) Retrieve and Verify: The top $N_{insp}=200$ retrieved images were manually verified and labeled as informative based on strict, visually identifiable morphological criteria corresponding to each phenophase. Images were classified as emergence if plants had fewer than four pairs of adult leaves; as flowering if open petals were visible; as seeding if seed pods were open with clearly visible seeds; and as senescence if green leaves were absent. 3) Export and Analyze: Observation dates were converted to Day-of-Year (DOY). Mean DOY values were compared among stages using ANOVA, followed by Tukey’s HSD to identify pairwise differences.
Results. INQUIRE-Search retrieved stage-specific phenological observations of common milkweed in southern Quebec, with verification success varying strongly by phenophase (Figure 9, top). Flowering and seeding yielded the largest datasets ($N_{ret}=169$, $Y=0.85$; $N_{ret}=161$, $Y=0.81$), while emergence and senescence produced smaller but still substantial datasets ($N_{ret}=45$, $Y=0.23$; $N_{ret}=52$, $Y=0.26$). As illustrated by the verified examples in Figure 9, these differences likely reflect both the visibility of diagnostic morphological cues and observer preferences for photographing conspicuous life stages.
@@
Figure 9: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for “flowering” phenological stage. (Bottom) Phenological stages occur at significant times throughout the year. (left) density of observations from INQUIRE-Search corresponding to different stages. (right) distribution of observations from INQUIRE-Search in different phenological stages visualized with results from ANOVA/Tukey test.
In contrast, iNaturalist’s controlled phenology annotations performed well only for visually salient stages. We applied iNaturalist’s controlled phenology annotations, mapping “No Flowers or Fruits” to emergence, “Flowers” to flowering, “Fruits or Seeds” to seeding, and “No Live Leaves” to senescence. After verifying up to the first 200 images per stage, baseline filtering yielded 5 emergence ($Y=0.03$), 196 flowering ($Y=0.98$), 52 seeding ($Y=0.26$), and 17 senescence images ($Y=0.09$). As a result, structured metadata produced an order-of-magnitude fewer usable records for early and late phenophases ($Y_{ratio}=7.7$ and $2.9$ for emergence and senescence respectively), whereas INQUIRE-Search consistently recovered informative examples across all four stages.
Using the verified INQUIRE-Search records, we observed a clear temporal progression across the growing season, with emergence occurring earliest, followed by flowering, seeding, and senescence (Figure 10, bottom), as expected. An ANOVA detected significant differences among stages in day-of-year distributions, and post-hoc Tukey tests indicated that all pairwise comparisons were significant (p < 0.05) except between emergence and flowering, which show substantial overlap in Figure 10. This overlap likely reflects both biological continuity between early growth and flowering and ambiguity in visual cues distinguishing these stages in images.
Dataset retrieved using INQUIRE-Search resolve fine-grained phenological structure at regional scales using community-science imagery. Emergence and senescence were more difficult to retrieve than flowering and seeding, reflecting subtle visual distinctions and observer biases toward photographing flowers and fruits (iNaturalist, 2022). Senescent observations distributed across much of the year, indicating delayed visibility that must be considered when interpreting phenological timing. Overall, these results demonstrate that INQUIRE-Search enables scalable, stage-specific phenological inference from opportunistic community-science imagery beyond what is feasible with traditional survey-based approaches.
3.5 CS5: Whale re-identification
Overview. This case study evaluates whether INQUIRE-Search can surface individually identifiable humpback whale flukes images from unstructured community-science imagery and link them to known individuals in existing ID catalogues.
Figure 10: (Left) Re-Identification of humpback whales. Unique Re-Identification of individuals is possible through careful examination of the fluke. (Right) iNaturalist humpback whale images with “fluke” description/tag filtering. Recent iNaturalist humpback whale images, even with the fluke filtering, are not identifiable due to the photographed angle or image quality.
Animal re-identification—the ability to recognize individual animals across space and time—is essential for population monitoring and movement ecology  (Krebs and others, 1989). For wide-ranging marine mammals such as humpback whales, these analyses rely on longitudinal photo-identification (photo-ID) datasets that link repeated sightings of the same individual across years and ocean basins using unique pigmentation and trailing-edge patterns on the ventral fluke, as shown in Figure 10  (Katona and Whitehead, 1981; Calambokidis et al., 2001; Martin et al., 1984; Howard et al., 2018). Although many fluke photographs exist in repositories such as iNaturalist, identifiable images are rare and dispersed across unstructured image collections, limiting the construction of re-ID datasets at scale. We evaluate whether INQUIRE-Search can address this bottleneck by systematically surfacing identifiable fluke images and linking them to known individuals in the HappyWhale dataset (Howard et al., 2018).
INQUIRE-Search workflow for CS3. 1) Query and Prioritize: We queried for “white underside of humpback whale fluke” with the humpback whale species filter. 2) Retrieve and Verify: Images were considered informative only if the ventral fluke was clearly visible, unobstructed, and in focus. Top $N_{insp}=200$ retrievals were inspected. 3) Export and Analyze: For analysis, we assess whether any iNaturalist retrievals corresponded to known individuals in the HappyWhale dataset. To accelerate matching, we cropped each image using Grounding DINO (Liu et al., 2024) and embedded both the iNaturalist and HappyWhale datasets using a multi-species re-identification model (Otarashvili et al., 2024). For each of the verified INQUIRE-Search images, we retrieved the top three closest HappyWhale candidates by embedding similarity and manually reviewed these candidates to confirm matches. Although not conducted by experts, humpback flukes possess distinct and easily recognizable pigmentation and trailing-edge patterns that allow reliable identification by non-experts.
Results. INQUIRE-Search efficiently surfaced high-quality humpback whale fluke photographs suitable for individual identification. Of the $N_{insp}=200$ retried images examined, $N_{ret}=153$ images contained unobstructed, properly oriented, sufficient quality flukes, resulting in a screening yield of $Y=0.77$
For comparison, we used keyword filtering as a baseline, since no controlled term exists for fluke images on iNaturalist. The keyword used for the search was “tail”, as it yielded the highest number of filtered retrievals. We used $N_{insp}=200$ results and and resulted in $N_{ret}=42$, resulting in =0.21=0.21. Thus, INQUIRE-Search produced far more usable re-ID observations ($Y_{ratio}=3.67$), reflected in substantially higher recall in the images most likely to support downstream matching.